OpenStudy (anonymous):

A small start-up company with 16 employees will select its 4 board members from among the employees. How many different groups of board members are possible?

OpenStudy (anonymous):

now this is just the same as the previous one. you need \[\dbinom{16}{4}\] do you know how to compute \[\dbinom{n}{k}\] ?

OpenStudy (anonymous):

i dont think the last one was right lol it was actually some really large number

OpenStudy (anonymous):

well then it was because 'itinerary' meant the order mattered. so it was \[10\times 9\times 8\times7=5040\]

OpenStudy (anonymous):

yes! so then i do 16*15*14*13?

OpenStudy (anonymous):

but in this problem order does not matter, just who is selected. so this time it really is \[\dbinom{16}{4}\]

OpenStudy (anonymous):

Yes, you start out with 16*15*14*13 but this case is different because in this case you do not care in which order the people were picked. so you have do divide by the number of ways you can permute the 4 people, which is 4*3*2

OpenStudy (anonymous):

divide first, multiply last. \[\dbinom{16}{4}=\frac{16 \times 15\times 14\times 13}{4 \times 3\times2}\]

OpenStudy (anonymous):

1820?

OpenStudy (anonymous):

\[=8\times 5\times 14\times 13\]

OpenStudy (anonymous):

oops sorry

OpenStudy (anonymous):

2*5*14*13=1820

OpenStudy (anonymous):

you are right. i canceled incorrectly. sorry.

OpenStudy (anonymous):

no worries i just multiplied and then divided im lucky ahah

OpenStudy (anonymous):

that always works with a calculator. but the answer is always a whole number (because the question is "how many") so you can always cancel the entire denominator.

OpenStudy (anonymous):

Oh I see, thank you very mch!

OpenStudy (anonymous):

welcome.